# A Survey on Vision-Language-Action Models for Embodied AI

### v4

@@ III-B2 Transformer-based Control Policies
RoboFlamingo [121] adapts the existing VLM, Flamingo [32, 149], to a robot policy by attaching an LSTM-based policy head to the VLM. This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
@@ III-B3 Control Policies for Multimodal Instructions
Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger.
@@
MOO [111] extends RT-1 to handle multimodal prompts. Leveraging the backbone of RT-1, MOO incorporates OWL-ViT to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
@@ III-B4 Control Policies with 3D Vision
We live in a 3D world, and it is intuitive that using 3D representations as visual inputs should provide richer information than 2D images. Point clouds are a popular choice for representing 3D inputs due to their straightforward derivation from RGBD inputs, as demonstrated by both DP3 and 3D Diffuser Actor. However, voxels have also been explored in various works. RoboUniView [124] shows improved performance by injecting 3D information into RoboFlamingo through a novel UVFormer module as its vision encoder, which provides 3D occupancy information from multi-perspective images. VER [151] also proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks.
PerAct [103] represents an advancement in both the observation and action spaces by leveraging 3D voxel representations. This approach offers a robust structural prior for action learning, enabling natural processing of multi-view observations and facilitating data augmentation in 6-DoF. In this framework, the input to the model comprises voxel maps reconstructed from RGBD images, while the output corresponds to the best voxel for guiding the gripper’s movement. By adopting this formulation, PerAct facilitates efficient task learning even with a few demonstrations. On the other hand, Act3D [104] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [105, 106] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
@@ III-B5 Diffusion-based Control Policies
Octo [134] introduces a transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer. Being among the first to utilize the Open X-Embodiment dataset [143], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
@@ III-B6 Diffusion-based Control Policies with 3D Vision
Several works have proposed combining 3D Vision with diffusion-based policies. DP3 [132] introduces 3D inputs to a diffusion policy, resulting in improved performance. Similarly, 3D Diffuser Actor [135] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy.
@@ III-B9 Large VLA
OpenVLA [50] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization.
@@
(a) FiLM
(b) Cross-attention
(c) Concatenation
Figure 5: Three common vision-language fusion methods of Transformer-based control policies. FiLM layers (Hadamard product $\odot$) are used in RT-1 models. Some utilize cross-attention to condition on instructions. Concatenation ($\oplus$) is the prevailing method in LVLAs.
Pros and Cons.
@@
Various VLA architectures explore different methods of fusing vision and language inputs, including cross-attention, FiLM, and concatenation, as illustrated in Figure 5. FiLM is used in RT-1 and thus its follow-up works inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [150].
@@
$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}})\end{split}$ (6)
@@
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right)$ (7)
@@
Table IV: High-level task planners. VL: vision-language fusion method.
Model V-Encoder Lang Model VL Control Policy Train Robotic Tasks
SayCan [10] - PaLM, FLAN - BC-Z, MT-Opt No Real-Navi&Mani (Everyday): office kitchen
@@
(SL)3 [155] - T5-small - seq2seq Yes Sim-Navi: ALFRED
EmbodiedGPT [156] EVA ViT-G/14 [157] LLaMA-7B Embodied-former MLP Yes Sim-Mani: Franka Kitchen, Meta-World
PaLM-E [11] ViT, OSRT [158] PaLM Concat Interactive Language, RT-1 Yes Sim: TAMP; Sim&Real-Mani: Lang.-Table; Real: SayCan setup
LEO [159] ConvNeXt [160], PointNet++ [161] Vicuna-7B Concat MLP Yes Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [57] CLIP, 3D-CLR [162], ConceptFusion [163] Flamingo, BLIP-2 Concat DD-PPO [164] Yes Sim-Navi: Habitat
ShapeLLM [165] ReCon++ LLaMA Concat - Yes 3D MM-Vet Benchmark
Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM Language CLIPort, MDETR, SayCan policies No Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [166] HLSM GPT-3 Language HLSM Yes Sim-Navi: ALFRED
LID [167] CNN GPT-2 Language & Embed - Yes Sim: VirtualHome
SMs [168] CLIP, ViLD GPT-3, RoBERTa [169] Language & Code CLIPort No Sim-Mani: pick, place
ConceptGraphs [170] SAM, CLIP, LLaVA [8] GPT-4 Code (JSON) API No Sim: AI2-THOR; Real (Spot Arm): pick, place
ProgPrompt [171] ViLD GPT-3 Code API No Sim: VirtualHome
ChatGPT for Robotics [172] YOLOv8 [173] ChatGPT Code API No Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [174] ViLD, MDETR GPT-3, Codex Code API No Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [175] CLIP ChatGPT Code MC-Controller, Steve-1 No Sim: Minecraft
@@ IV Task Planners
A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as a language instruction to low-level control policies $\pi_{\theta}$:
@@
This process is often referred to as task decomposition and is also related to the task and motion planning (TAMP) setting. Ideally, task plans should also incorporate optimal scheduling for these subtasks. When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4. Table IV outlines important details about task planners.
@@ IV-A Monolithic Task Planners
IV-A1 Grounded Task Planners
@@
IV-A2 End-to-end Task Planners
Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning. Clearly, they are also monolithic models.
IV-A3 End-to-end Task Planners with 3D Vision
LEO [159] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO demonstrates proficiency not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
3D-LLM [57] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog, and navigation. The 3D features can come in diverse forms, including point cloud, gradSLAM, and neural voxel field. MultiPLY [179] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
@@ IV-A3 End-to-end Task Planners with 3D Vision
IV-B Modular Task Planners
Finetuning end-to-end models on embodied data can be expensive, and there are approaches that adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners.
@@ IV-B1 Language-based Task Planners
Table V: Recent datasets for robot learning. Following the definition in RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulators rather than in the real world. Some datasets are continuously updated; this table is based on the information from their original papers. Table adapted from [180, 146].
Name Collection Visual Instruction Embodiment Skills Tasks Scenes Objects Episodes
MIME [181] Human Teleoperation RGB, D Demo Baxter 12 20 1 - 8.3K
RoboTurk [182] $\ast$ Human Teleoperation RGB - Sawyer 2 - 1 - 2.1K
RoboNet [183] Scripted RGB Goal image (7 robots) - - 10 - 162K
MT-Opt [184] Scripted RGB Lang (7 robots) 2 12 1 - 800K
BC-Z [113] Human Teleoperation RGB Lang, Demo Everyday 3 100 1 - 25.9K
RT-1-Kitchen [114] Human Teleoperation RGB Lang Everyday 12 700+ 2 16 130K
MOO [111] Human Teleoperation RGB Multi-modal Everyday 5 - - 106 59.1K
VIMA [150] $\ast$ Scripted RGB Multi-modal UR5 17 - 1 29 650K
RoboSet [185] Human, Scripted RGB, D Lang Franka Panda 12 38 11 - 98.5K
BridgeData V2 [186] Human, Scripted RGB, D Lang WidowX 250 13 - 24 100+ 60.1K
RH20T [180] Human Teleoperation RGB, D Lang (4 robots) 42 147 7 - 110K+
DROID [146] Human Teleoperation RGB, D Lang Franka Panda 86 - 564 - 76K
OXE [143] Aggregated Datasets RGB, D Lang (22 robots) 527 160,266 311 - 1M+
@@ IV-B2 Code-based Task Planners
Pros and Cons. Monolithic task planners that utilize grounded task planning focus on generating executable plans. End-to-end models like PaLM-E share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance. However, the training costs of such large models can be a significant concern. In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs. Language-based task planners offer the advantage of seamless integration of LLMs and VLMs, as they are designed to operate in the natural language space. However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies. On the other hand, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability. Nevertheless, their performance can be constrained by the programming capabilities of existing models.
@@ V Datasets and Benchmarks
Table VI: Simulators and simulated benchmarks. Control: Continuous control tasks. D, S, A, N: Depth, segmentation, audio, normals. Force: simulated contact force between end-effector and item. PD: Pre-defined. Table adapted from [188, 189].
@@
Gibson [190] 572/- - - Pybullet Navi RGB, D, N, S - - Navi only -
iGibson [191, 192, 189] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Versions: iGibson 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [188], BEHAVIOR-1K [193]
SAPIEN [194] - 2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmarks: SIMPLER [195]
@@
TDW [201] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids -
@@
CALVIN [203] 4/- 7/5 - Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
@@
Habitat [205, 206] Matterport + Gibson Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: Rearrangement [207], Habitat 2.0 VC-1, PACT; OVMM [208]
ALFRED [209] -/120 84/84 - Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
DMC [210] 1/- 4/4 Code MuJoCo Control RGB, D Force - Continuous RL VC-1, SMART
OpenAI Gym [211] 1/- 4/4 Code MuJoCo Control RGB Force - Single agent RL environments -
Genesis [212] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation -
V-A Real-world Robot Datasets & Benchmarks
V-B Simulators, Simulated Robot Datasets & Benchmarks

### v5

@@ III-B2 Transformer-based Control Policies
RoboFlamingo [99] adapts the existing Flamingo VLM to a robot policy by attaching an LSTM-based policy head to the VLM. This demonstrates that pretrained VLMs can be effectively transferred to language-conditioned robotic manipulation tasks.
(a) FiLM
(b) Cross-attention
(c) Concatenation
(d) Quantization
(e) Tool Use
Figure 5: Representative architectures of VLA models. $\odot$: Hadamard product. $\oplus$: Concatenation.
@@ III-B3 Control Policies for Multimodal Instructions
Multimodal instruction enables new ways to specify tasks, such as through demonstrations, by naming novel objects, or by pointing with a finger or mouse click.
@@
MOO [93] extends RT-1 to handle multimodal prompts. Leveraging the backbone of RT-1, it incorporates OWL-ViT to encode images within the prompt. By expanding the RT-1 dataset with new objects and additional prompt images, MOO enhances the generalization capabilities of RT-1. This extension also facilitates new methods of specifying target objects, such as pointing with a finger, and clicking on graphical user interfaces.
@@ III-B4 Control Policies with 3D Vision
In our 3D world, it is reasonable to assume that 3D vision provides richer information than 2D images.
Point clouds are a common representation for 3D visual inputs due to their straightforward derivation from RGB-D inputs, as demonstrated by both DP3 [105] and 3D Diffuser Actor [108]. Voxels have likewise been extensively studied. VER [127] proposes voxelizing multi-view images into 3D cells in a coarse-to-fine manner, enhancing performance in vision-language navigation tasks. PerAct [87] facilitates efficient task learning with only a few demonstrations by leveraging voxel representations in both the observation and action spaces. The input comprises voxel maps reconstructed from multi-view RGB-D images, while the output corresponds to the best voxel for guiding the gripper’s movement. RoboUniView [100] improves performance by injecting 3D information from multi-perspective images through a novel UVFormer vision encoder, which is pretrained on a 3D occupancy task.
In contrast, Act3D [88] introduces a continuous resolution 3D feature field, with adaptive resolutions based on current task, addressing the computational cost of voxelization. RVT, RVT-2 [89, 90] propose re-rendering images from virtual views of the scene point cloud and using these images as inputs, rather than directly relying on 3D inputs.
@@ III-B5 Diffusion-based Control Policies
Octo [107] introduces a Transformer-based diffusion policy characterized by a modular open-framework design, allowing for flexible connections from different task definition encoders, observation encoders, and action decoders to the Octo Transformer. Being among the first to utilize the OXE dataset [112], Octo demonstrates positive transfer and generalizability across diverse robots and tasks.
@@ III-B6 Diffusion-based Control Policies with 3D Vision
Several works have proposed combining 3D Vision with diffusion-based policies. DP3 [105] introduces 3D inputs to a diffusion policy, resulting in improved performance. Similarly, 3D Diffuser Actor [108] shares the core idea of DP3 but differs in model architecture by combining Act3D with Diffusion Policy. 3D-MoE [130] explores an efficient mixture-of-experts architecture with DiT-based action diffusion using a rectified flow scheduler.
@@ III-B9 Large VLA
OpenVLA [37] was later developed as an open-source counterpart to RT-2-X. They additionally explored efficient fine-tuning methods, including LoRA and model quantization. OpenVLA-OFT [113] applies an Optimized Fine-Tuning (OFT) recipe for improved efficiency and performance. TraceVLA [114] finetunes OpenVLA to enable visual trace prompting, enhancing spatial-temporal awareness.
@@
RoboMamba [116] replaces the costly Transformer with a Mamba state space model that has linear inference complexity, achieving efficient robotic reasoning and action capabilities.
SpatialVLA [117] introduces Ego3D Position Encoding for injecting 3D information and represents actions with Adaptive Action Grids for enhanced generalizability and robustness.
LAPA [121] devises the first unsupervised pretraining method for VLAs based on latent actions [69]. This approach employs a three-stage process to learn from internet-scale unlabeled videos. First, a VQ-VAE framework is pretrained to extract quantized latent actions between image frames. Next, a VLA model is pretrained to predict these latent actions. Finally, the model is finetuned using only a small robotic dataset to map the latent actions to actual robot actions.
The integration of large VLAs with diffusion has also been gaining popularity. TinyVLA [118] leverages the Diffusion Policy, while CogACT [119] utilizes a DiT action diffusion module. DexVLA [20] proposes embodied curriculum learning to progressively train a diffusion action expert, incorporating sub-step reasoning for decomposing long-horizon tasks. HybridVLA [120] integrates diffusion with the autoregression paradigm to fully leverage VLMs’ reasoning capabilities.
Visual autoregressive modeling (VAR) [133] with quantized visual tokens demonstrates improved performance over diffusion models in image generation. This suggests that the three modalities of VLAs can be unified under the autoregressive paradigm [134]. WorldVLA [122] and UniVLA [123] advance this direction by integrating VLAs with world models. It quantizes multimodal data into discrete tokens, forming a shared vocabulary of quantized multimodal tokens. Consequently, all modalities can be modeled autoregressively, enabling not only action and text generation but also image generation, thereby constituting a world model.
A recent trend in LLMs is equipping them with tool-use capabilities by generating code that calls tools via APIs [135]. Instruct2Act [102] follows this paradigm by integrating vision and action tools, enabling LLMs to perform robotic tasks.
Strengths and Limitations.
@@
Representative VLA architectures are illustrated in Figure 5. FiLM is used in RT-1 and thus its follow-up models inherit this mechanism. While cross-attention may offer superior performance with smaller model sizes, concatenation is simpler to implement and can achieve comparable results with larger models [126]. Quantization unifies multimodal tokens into a shared vocabulary, thus enabling integration with world models. The tool-use paradigm of LLMs can also be applied to robotic tasks.
@@
$\begin{split}\mathcal{L}_{SE(2)}&=\operatorname{CE}(a_{\text{pick}},\hat{a}_{\text{pick}})+\operatorname{CE}(a_{\text{place}},\hat{a}_{\text{place}}).\end{split}$ (6)
@@
$\mathcal{L}_{\text{DDPM}}=\operatorname{MSE}\left(\varepsilon^{k},\varepsilon_{\theta}(a_{t}+\varepsilon^{k},k)\right),$ (7)
@@
Table IV: High-level task planners.
Model Vision
Model
Language Model / VLM Low-level Control Policy Architecture Require Training Environments, Embodiments, Tasks, and Skills
PaLM-E [11] ViT, OSRT PaLM Interactive Language, RT-1 Concat Sim: TAMP; Sim&Real-Mani: Language-Table; Real: SayCan setup
EmbodiedGPT [136] EVA ViT-G/14 LLaMA-7B MLP Xattn Sim-Mani: Franka Kitchen, Meta-World
LEO [137] ConvNeXt, PointNet++ Vicuna-7B MLP Concat Sim-Mani: CLIPort task subset; Sim-Navi: Habitat-web
3D-LLM [44] CLIP, 3D-CLR, ConceptFusion Flamingo, BLIP-2 $D$-PPO Concat Sim-Navi: Habitat
ShapeLLM [138] ReCon++ LLaMA   Concat 3D MM-Vet Benchmark
SayCan [10]   PaLM, FLAN BC-Z, MT-Opt   Real-Navi&Mani (EDR): office kitchen
@@
(SL)3 [140]   T5-small seq2seq   Sim-Navi: ALFRED
Inner Monologue [9] MDETR, CLIP InstructGPT, PaLM CLIPort, heuristics, SayCan policies Language Sim&Real-Mani: Tabletop Rearrangement; Real-Navi&Mani: Kitchen Mobile Mani.
LLM-Planner [141] HLSM object detector GPT-3 HLSM Language Sim-Navi: ALFRED
LID [142] CNN GPT-2   Lang., Embed Sim: VirtualHome
SMs [143] CLIP, ViLD GPT-3, RoBERTa CLIPort Lang., Code Sim-Mani: pick, place
ProgPrompt [144] ViLD GPT-3 API Code Sim: VirtualHome
ChatGPT for Robotics [145] YOLOv8 ChatGPT API Code Real-Navi: drone flight; Sim-Navi: AirSim, Habitat
CaP [146] ViLD, MDETR GPT-3, Codex API Code Sim-Mani & Sim-Navi: pick, place, etc.
DEPS [147] CLIP ChatGPT MC-Controller, Steve-1 Code Sim: Minecraft
ConceptGraphs [148] SAM, CLIP, LLaVA GPT-4 API Code (JSON) Sim: AI2-THOR; Real (Spot Arm): pick-place
@@
Scaling Law
Similar to LLMs, scaling laws have also been observed in robotics [149], revealing the importance of model size, data quality, and diversity in environments and objects. Further research in this area can guide the development of VLAs with improved in-the-wild generalization capabilities.
@@ IV Task Planners
A high-level task planner $\pi_{\phi}$ aims to divide a complex task $\ell$ into a sequence of subtasks $p_{i}$ (i.e., task plan), each serving as an instruction to low-level control policies $\pi_{\theta}$:
@@
This process is sometimes referred to as task or subgoal decomposition and is closely related to task and motion planning (TAMP) and embodied decision making. When equipped with task planners, VLAs can complete more complex, long-horizon tasks, as illustrated in Figure 4. Details are summarized in Table IV. Ideally, task plans should also incorporate optimal scheduling for these subtasks.
@@ IV-A Monolithic Task Planners
IV-A1 End-to-end Task Planners
Similar to LVLAs, task planners can be implemented as end-to-end multimodal LLMs, leveraging their Internet-scale knowledge for task planning.
IV-A2 End-to-end Task Planners with 3D Vision
LEO [137] uses a two-stage training strategy to integrate a point cloud encoder with an LLM: the first stage focuses on 3D vision-language alignment, followed by the second stage, which involves 3D vision-language-action instruction tuning. LEO performs well not only in 3D question-answering tasks but also in manipulation, navigation, and task planning.
3D-LLM [44] injects 3D information into LLMs and empowers them to perform 3D tasks, such as 3D-assisted dialog and navigation, using features from point cloud, gradSLAM, and neural voxel field. MultiPLY [150] is an object-centric embodied LLM that incorporates even more modalities, including audio, tactile, and thermal.
IV-A3 Grounded Task Planners
@@
IV-B Modular Task Planners
Finetuning end-to-end models on embodied data can be expensive, and some approaches adopt a modular design by assembling off-the-shelf LLMs and VLMs into task planners. These approaches can also be viewed as following the tool-use architecture [135].
@@ IV-B2 Code-based Task Planners
Strengths and Limitations.
Monolithic task planners that utilize grounded task planning focus on generating executable plans. End-to-end models share an architecture similar to most LVLAs and can be finetuned on specialized embodied data to achieve better performance. However, the training costs of such large models can be a concern. In contrast, modular task planners are more readily deployable because they leverage off-the-shelf LLMs and VLMs. Language-based task planners offer the advantage of seamless integration of LLMs and VLMs, as they are designed to operate in the natural language space. However, they often require extra steps to align the generated task plans with language instructions that are admissible to low-level control policies. Conversely, while code-based task planners may require manually wrapping VLMs and control policies in APIs and preparing clear documentation in advance, they enable code debugging and provide greater controllability. Nevertheless, their performance can be constrained by the programming capabilities of existing models.
@@ V Datasets and Benchmarks
V-A Real-world Robot Datasets & Benchmarks
V-B Simulators, Simulated Robot Datasets & Benchmarks
Table V: Embodied Datasets. Per RT-X, skills correspond to verbs, and tasks are different combinations of verbs and objects. $\ast$ Dataset collected in simulation rather than the real world. Some datasets are continually updated, and we include only the original version. Table adapted from [153, 154].
Dataset Skills Tasks Scenes Episodes Collection Obs. Instruction Robots
MIME [155] 12 20 1 8.3K Human teleop. RGBD Demo Baxter
[…]
RoboNet [157]     10 162K Script RGB Goal image (7 robots)
MT-Opt [158] 2 12 1 800K Script RGB Lang (7 robots)
BC-Z [79] 3 100 1 25.9K Human teleop. RGB Lang, demo EDR
Fractal [94] 12 700+ 2 130K Human teleop. RGB Lang EDR
MOO [93] 5     59.1K Human teleop. RGB Multimodal EDR
[…]
RoboSet [159] 12 38 11 98.5K Human, script RGBD Lang Franka
BridgeV2 [160] 13   24 60.1K Human, script RGBD Lang WidowX
RH20T [153] 42 147 7 110K+ Human teleop. RGBD Lang (4 robots)
DROID [154] 86   564 76K Human teleop. RGBD Lang Franka
OXE [112] 527 160K 311 1M+ Aggregate data RGBD Lang (22 robots)
Table VI: Simulators and simulated benchmarks. Control: continuous control tasks. D, S, A, N: depth, segmentation, audio, normal. Force: simulated contact force between end-effector and item. PD: pre-defined. Table adapted from [161, 162].
@@ V-B Simulators, Simulated Robot Datasets & Benchmarks
Gibson [163] 572/-     Pybullet Navi RGB, D, N, S     Navi only
iGibson [164, 165, 162] 15/108 152/5 Mouse, VR Pybullet Navi, Mani RGB, D, S, N, Flow, LiDAR Force TurtleBot v2, LoCoBot, etc. VR, Continuous Extended States. Versions: 0.5, 1.0, 2.0 Benchmarks: BEHAVIOR-100 [161], BEHAVIOR-1K [166]
SAPIEN [167]   2346/- Code PhysX Navi, Mani RGB, D, S Force Franka Articulation, Ray Tracing VoxPoser. Benchmark: SIMPLER [168]
@@
TDW [174] 15/120 112/50 VR Unity, Flex Navi, Mani RGB, D, S, A Force Fetch, Sawyer, Baxter Audio, Fluids
@@
CALVIN [177] 4/- 7/5   Pybullet Mani RGB, D Force Franka Long-horizon Lang-cond tasks GR-1, HULC, RoboFlamingo
@@
Habitat [179, 180] (Matterport + Gibson) Mouse Bullet Navi RGB, D, S, A Force Fetch, Franka, AlienGO Fast, Navi only. Versions: 1.0, 2.0, Rearrangement [181] VC-1, PACT; Benchmark: OVMM [182]
ALFRED [183] -/120 84/84   Unity Navi, Mani RGB, D, S PD Human Diverse long-horizon tasks (SL)3, LLM-Planner
DMC [184] 1/- 4/4 Code MuJoCo Control RGB, D Force   Continuous RL VC-1, SMART
OpenAI Gym [185] 1/- 4/4 Code MuJoCo Control RGB Force   Single agent RL environments
Genesis [186] (Rigid, deformable, liquid, etc.) Code (Propri-etary) Navi, Mani RGB, D, S, N Force Franka, Unitree, etc. High-speed comprehensive physics simulation